If it is known that sin a = 4/5.
pi/2 < a < pi
and that sin b = -2sqrt(5)/5
pi < b < 3pi/2 find the exact value of cos(a + b)
I am completely lost on this problem, any and all help would be greatly appreciated

Answer by stanbon(75887)

If it is known that sin a = 4/5.
pi/2 < a < pi ::: QII where x <0 and y > 0
Since sin = y/r, y = 4 and r = 5
Then x = -sqrt[25-16] = -3
---
So cos(a) = x/r = -3/5
---------------------------------------------------
and that sin b = -2sqrt(5)/5

pi < b < 3pi/2 ::: QIII where x < 0 and y < 0
Since sin = y/r, y = -2sqrt(5) and r = 5
Then x = -sqrt[25-20] = -sqrt(5)
----
so cos(b) = -sqrt(5)/5
---------------------
find the exact value of cos(a + b)
cos(a+b) = cos(a)cos(b)-sin(a)sin(b)
---
= (-3/5)(-sqrt(5)/5) - (4/5)(-2sqrt(5)/5)
===========================================
Cheers,
Stan H.
==================